OpenStudy (anonymous):

I am thinking of a number x, on condition 0 < x < 10. If I double it, then add 6, then divide the result by 2, then subtract the number I am thinking of and I get 3. What number was I thinking of? @satellite73 @ganeshie8 @AravindG @Directrix @mathslover @robtobey @primeralph @Luigi0210 @shamil98 @ranga @charlotte123 @Zale101 @kc_kennylau @GoldPhenoix @Compassionate @jagatuba @ChmE @KattGirl14 @helpme1.2 @Yacoub1993 @divu.mkr @nikato @Kristen17 @raffle_snaffle @ArkGoLucky @ikram002p @armine @petiteme @demonguy @youngamerican @creeksider @rodrigorodriguez13 @Fellowroot @nannaberry @kittiwitti1

OpenStudy (anonymous):

\[ \frac{1}{2} (2 x+6)-x=3 \\ x +3 - x =3\\ 0 x =0 \] x can be any number

OpenStudy (anonymous):

between 0 and 10

OpenStudy (goldphenoix):

If so, then check your answer*

OpenStudy (goldphenoix):

My bad, I didn't see the "add 6". Sorry for doubting you, @eliassaab

OpenStudy (kittiwitti1):

No tagging please (: PMs are fine but mass tagging can end up getting you reported ^^